OpenStudy (anonymous):

How many sulfur atoms are there in 4.55mol of sulfur? Express your answer using three significant figures.

OpenStudy (abb0t):

Are you familiar with avogadro's constant? 6.022 \(\times\) 10\(^{23}\) \(\frac{mol}{atom}\)

OpenStudy (anonymous):

No, if you don't mind can you explain.

OpenStudy (abb0t):

You will be using dimensional analysis here. You're given: 4.55 mol of S \(\times \frac{6.022 \times 10^{23} atoms}{mol}\)